OpenStudy (anonymous):

Find the GCF (Greatest Common Factor) of this pair of monomials: 15x^4 and 35x^2

OpenStudy (jdoe0001):

hint: \(15x^4 \qquad 35x^2 \implies \begin{cases} 15x^4\to 3\cdot {\color{brown}{ 5}}\cdot x\cdot x\cdot {\color{brown}{ x\cdot x }} \\ \quad \\ 35x^2\to {\color{brown}{ 5}}\cdot 7\cdot {\color{brown}{ x\cdot x }} \end{cases}\)

OpenStudy (skullpatrol):